# Block encoding of matrix product operators

### v2

@@ VI Applications
This section demonstrates how our block encoding technique can be applied to specific Hamiltonians. As a first example, we consider the Ising and Heisenberg Hamiltonians, which have a relatively simple MPO representation, and a slightly modified version of the XY model with exponentially decaying potential.
Then, we consider the case where a sum of Pauli strings can be rearranged in a tensor product of sums of local Pauli operators. For this system, we also perform an eigenstate filtering procedure through an entire QET circuit.
@@ VI.1 Ising model with transverse field
with $Z_{\ell}$ and $X_{\ell}$ Pauli operators acting on the $\ell$-th site. Then, a suitable MPO decomposition can be achieved by defining the internal $A^{(\ell)}$ tensors as [20, 7]:
$A^{(\ell)}=\begin{pmatrix}I_{\ell}&0&0\\ Z_{\ell}&0&0\\ gX_{\ell}&JZ_{\ell}&I_{\ell}\end{pmatrix}$ (23a)
and by “closing” the tensor chain with the following initial and final vectors of operators, respectively:
$\tilde{A}^{(1)}$ $=\begin{pmatrix}gX_{1}&JZ_{1}&I_{1}\end{pmatrix},$ (23b)
$\tilde{A}^{(L)}$ $=\begin{pmatrix}I_{L}\\ Z_{L}\\ gX_{L}\end{pmatrix}.$ (23c)
Every local operator $O_{\ell}$ acts on the $\ell$-th lattice site, and $I$ is the $2\times 2$ identity matrix. Such a representation can be easily found through a finite state automata picture of the Hamiltonian, as shown in [7, 16].
@@
$\begin{pmatrix}gX_{1}&JZ_{1}&I\end{pmatrix}$ $=\begin{pmatrix}0&0&I\end{pmatrix}\cdot A^{(1)},$ (24a)
$\begin{pmatrix}I\\ Z_{L}\\ gX_{L}\end{pmatrix}$ $=A^{(L)}\cdot\begin{pmatrix}I\\ 0\\ 0\end{pmatrix}$ (24b)
@@
$\mathcal{H}_{\text{Ising}}=\begin{pmatrix}0&0&I\end{pmatrix}\left[\;\prod_{\ell=1}^{L}\begin{pmatrix}I&0&0\\ Z_{\ell}&0&0\\ gX_{\ell}&JZ_{\ell}&I\end{pmatrix}\;\right]\begin{pmatrix}I\\ 0\\ 0\end{pmatrix}.$ (25)
We immediately notice that since the Pauli operators act on single qubits, and have dimension $2\times 2$, each matrix has dimension $6\times 6$ and therefore needs at least 3 qubits to be represented. As a consequence, all the matrices and vectors in Eq. (25) gain a new row and/or column in the following way:
@@
$A^{(\ell)}$ $\longrightarrow\begin{pmatrix}I&0&0&0\\ Z_{\ell}&0&0&0\\ gX_{\ell}&JZ_{\ell}&I&0\\ 0&0&0&I\end{pmatrix}.$ (26c)
@@
Note that to construct a block encoding of $A^{(\ell)}$, we have to normalize it by a factor $N\geq\norm{\mathcal{M}\left[A^{(\ell)}\right]}$. This has the same effect as dividing the Hamiltonian parameters $J$ and $g$ by $N^{L}$:
@@
$A^{(\ell)}$ $\longrightarrow\begin{pmatrix}I&0&0&0\\ Z_{\ell}&0&0&0\\ gX_{\ell}&JZ_{\ell}&I&0\\ 0&0&0&\zeta^{1/L}I\end{pmatrix}$ (28c)
@@
$J$ $\longrightarrow\frac{J}{2N^{L}},$ (29ada)
$g$ $\longrightarrow\frac{g}{2N^{L}}.$ (29adb)
@@ VI.2 Heisenberg model
Next, we discuss the Heisenberg model governed by the Hamiltonian
$\begin{split}\mathcal{H}_{\text{Heis}}&=\sum_{\ell=1}^{L-1}\left(J_{X}X_{\ell}X_{\ell+1}+J_{Y}Y_{\ell}Y_{\ell+1}+J_{Z}Z_{\ell}Z_{\ell+1}\right)\\ &+\sum_{\ell=1}^{L}\left(g_{X}X_{\ell}+g_{Y}Y_{\ell}+g_{Z}Z_{\ell}\right).\end{split}$ (29ae)
@@
$A^{(\ell)}=\begin{pmatrix}I&0&0&0&0\\ X_{\ell}&0&0&0&0\\ Y_{\ell}&0&0&0&0\\ Z_{\ell}&0&0&0&0\\ g_{X}X_{\ell}\\ +g_{Y}Y_{\ell}+g_{Z}Z_{\ell}&J_{X}X_{\ell}&J_{Y}Y_{\ell}&J_{Z}Z_{\ell}&I\\ \end{pmatrix}.$ (29af)
The matrix $A^{({\ell})}$ in Eq. (29af) has dimension $10\times 10$ and therefore needs to be expanded into a $16\times 16$ matrix (analogous to the Ising MPO model) to be represented by $4$ qubits. Moreover, another dilation ancillary is added to build $U_{A^{(\ell)}}$ through a QR decomposition.
@@ VI.3 Exponentially decaying XY model
$\mathcal{H}_{\text{XY}_{\alpha}}=\sum_{\ell_{1}=1}^{L}\sum_{\ell_{2}=\ell_{1}+1}^{L}e^{-\alpha|\ell_{1}-\ell_{2}|}\left(J_{X}X_{\ell_{1}}X_{\ell_{2}}+J_{Y}Y_{\ell_{1}}Y_{\ell_{2}}\right)$ (29ag)
with $\alpha>0$.
If we wanted to represent such a Hamiltonian with a linear combination of unitaries, we would need to consider $M=\mathcal{O}(L^{2})$ terms, which results in a computational complexity of $\mathcal{O}(L^{2}\cdot\log^{2}(L)$ and $\mathcal{O}(\log(L))$ ancillaries.
@@
$A^{(\ell)}=\begin{pmatrix}I&0&0&0\\ e^{-\alpha}X_{\ell}&e^{-\alpha}I&0&0\\ e^{-\alpha}Y_{\ell}&0&e^{-\alpha}I&0\\ 0&J_{X}X_{\ell}&J_{Y}Y_{\ell}&I\\ \end{pmatrix}.$ (29ah)
Our MPO block encoding technique would then be able to represent this Hamiltonian with a number of one- and two-qubit scaling only linearly with the system size $L$. The number of ancillaries would also scale as $\mathcal{O}(L)$ as well as the cost of the signal processing, giving a total cost for a full QET circuit of only $\mathcal{O}(d\times L)$, where $d$ is the degree of the polynomial. It’s also worth noticing that we don’t need to consider exponentially decaying coefficients, as the factor $e^{-\alpha}$ is the same in every matrix $A^{(\ell)}$.
We have then found a first example of a practical application where our technique can constitute a significant advantage compared to the previous state-of-the-art method.
VI.4 Tensor product of sum of Paulis
@@
$\mathcal{H}_{\text{TP}}=\bigotimes_{\ell=1}^{L}\left(\alpha_{\ell}I+\beta_{\ell}X_{\ell}+\gamma_{\ell}Y_{\ell}+\delta_{\ell}Z_{\ell}\right).$ (29ai)
@@
$A^{(\ell)}=\left(\alpha_{\ell}I+\beta_{\ell}X_{\ell}+\gamma_{\ell}Y_{\ell}+\delta_{\ell}Z_{\ell}\right)$ (29aj)
@@
$A^{(\ell)}\longrightarrow\begin{pmatrix}\alpha_{\ell}I+\beta_{\ell}X_{\ell}+\gamma_{\ell}Y_{\ell}+\delta_{\ell}Z_{\ell}&0\\ 0&\zeta^{1/L}I\end{pmatrix}$ (29ak)
@@
$\bigotimes_{\ell=1}^{L}\left(\alpha_{\ell}I+\beta_{\ell}X_{\ell}+\gamma_{\ell}Y_{\ell}+\delta_{\ell}Z_{\ell}\right)+\zeta I=\\ \begin{pmatrix}I&I\end{pmatrix}\left[\prod_{\ell=1}^{L}\begin{pmatrix}\alpha_{\ell}I+\beta_{\ell}X_{\ell}\\ +\gamma_{\ell}Y_{\ell}+\delta_{\ell}Z_{\ell}&0\\ 0&\zeta^{1/L}I\end{pmatrix}\right]\begin{pmatrix}I\\ I\end{pmatrix},$ (29an)
@@
VI.5 Eigenstate filtering
Figure 7: 7a Original eigenvalues for tensor product of Pauli sums with $L=3$, $\alpha=\{0.7,1.2,-0.3\}$, $\beta=\{-1,0.4,0.5\}$, $\gamma=\{0,0.3,0.5\}$, $\delta=\{0.1,0,1.2\}$, $\zeta=1.7$ and $N=1.72$. 7b Filtering function (in pink) plus QET-transformed eigenvalues. Phases were calculated via the pyqsp package with degree $d=30$ and threshold $\Delta_{\lambda}=0.1$.
Finally, we test our block encoding algorithm, as well as our signal processing technique, by applying an eigenstate filtering function to the Hamiltonian defined in (29ai). In particular, we want to perform the following transformation over $\mathcal{H}$ [22]:
$f_{d}(x,\Delta_{\lambda})=\frac{T_{d}\left(-1+2\frac{x^{2}-\Delta_{\lambda}^{2}}{1-\Delta_{\lambda}^{2}}\right)}{T_{d}\left(-1+2\frac{-\Delta_{\lambda}^{2}}{1-\Delta_{\lambda}^{2}}\right)}$ (29ao)
@@ VI.5 Eigenstate filtering
$\text{poly}(\mathcal{H})=\frac{1}{2}\left(f_{\text{QET}}(\mathcal{H})+f_{\text{QET}}(\mathcal{H})^{\dagger}\right).$ (29ap)
@@ VII Conclusion
Our block encoding method is generally applicable since typical Hamiltonians of interest can be represented as matrix product operators. Moreover, our block encoding technique’s computational cost grows only linearly with the system size and quadratically with the virtual bond dimension, offering a potential advantage compared to LCU in terms of one- and two-qubit gate count. In section VI we have shown one specific example of a computational advantage in the case of an exponentially decaying XY model. We remark that our technique uses more ancillary qubits than LCU.
The signal processing operator can have different implementations, as we have shown in section  IV, and it incurs a cost growing linearly or quadratically with the number of ancillaries, depending on the chosen implementation and the presence of one extra auxiliary qubit [18, 15].
As previously mentioned, each matrix encoded into a larger unitary must have a bounded norm. Encoding directly $\mathcal{H}$ in a larger unitary, or using LCU, requires $\norm{\mathcal{H}}\leq 1$. In our case, we need to normalize each tensor $A^{(\ell)}$ individually, causing an equivalent rescaling of the Hamiltonian by a factor $\prod_{\ell}^{L}N_{\ell}$. The required renormalization is a general issue of block encoding and is not specific to our MPO approach. We remark that the spectral norms of the individual matrices $\mathcal{M}[A^{(\ell)}]$ are typically easier to calculate than $\norm{\mathcal{H}}$. However, $\prod_{\ell}^{L}N_{\ell}$ can be larger than $\norm{\mathcal{H}}$. Note that the final success probability for a QET circuit is not directly impacted by the normalization factor, which appears instead in the polynomial transformation argument, see Eq. (IV). Still, an exponential suppression of the eigenvalues might bring some practical disadvantages. We are leaving the question of how to minimize the norms $\|\mathcal{M}[A^{(\ell)}]\|$ for future work.

### v3

@@ VI Applications
This section demonstrates how our block encoding technique can be applied to specific Hamiltonians. As a first example, we consider the Ising and Heisenberg Hamiltonians, which have a relatively simple MPO representation and a slightly modified version of the XY model with exponentially decaying potential, where we notice a computational speed-up compared to the use of LCU. Then, we briefly comment on the case of a spinless one-dimensional Fermi-Hubbard Hamiltonian as a simple example of a fermionic system. Finally, we consider the case where a sum of Pauli strings can be rearranged in a tensor product of sums of local Pauli operators. We also perform an eigenstate filtering procedure for this system through a complete QET circuit.
In appendix B, we report a numerical analysis regarding the post-selection success probabilities for the physical systems considered in this section, comparing both MPO and LCU block encodings.
It’s worth mentioning that even if each of these examples shows a constant virtual bond dimension $\chi$, this is not the most general case. As a matter of fact, $\chi$’s scaling w.r.t the system size $L$ can vary based on the problem considered. Still, the analysis in section V can be applied in every circumstance.
@@ VI.1 Ising model with transverse field
with $Z_{\ell}$ and $X_{\ell}$ Pauli operators acting on the $\ell$-th site. A suitable MPO decomposition can be achieved by defining the internal $A^{(\ell)}$ tensors as [20, 7]:
$A^{(\ell)}=\begin{pmatrix}I&0&0\\ \sqrt{|J|}Z_{\ell}&0&0\\ gX_{\ell}&\text{sign}(J)\sqrt{|J|}Z_{\ell}&I\end{pmatrix}$ (26a)
where this MPO form was specifically chosen to ensure that $\left\|\mathcal{M}\left[A^{(\ell)}\right]\right\|\rightarrow 1$ for vanishing $g$ and $J$.
Then, we can “close” the tensor chain with the following initial and final vectors of operators, respectively:
$\tilde{A}^{(1)}$ $=\begin{pmatrix}gX_{1}&\text{sign}(J)\sqrt{|J|}Z_{1}&I\end{pmatrix},$ (26b)
$\tilde{A}^{(L)}$ $=\begin{pmatrix}I\\ \sqrt{|J|}Z_{L}\\ gX_{L}\end{pmatrix}.$ (26c)
Every local operator $O_{\ell}$ acts on the $\ell$-th lattice site, and $I$ is the $2^{L}\times 2^{L}$ identity matrix. Such a representation can be easily found through a finite state automata picture of the Hamiltonian, as shown in [7, 16].
@@
$\begin{pmatrix}gX_{1}&\text{sign}(J)\sqrt{|J|}Z_{1}&I\end{pmatrix}$ $=\begin{pmatrix}0&0&I\end{pmatrix}\cdot A^{(1)},$ (27a)
$\begin{pmatrix}I\\ \sqrt{|J|}Z_{L}\\ gX_{L}\end{pmatrix}$ $=A^{(L)}\cdot\begin{pmatrix}I\\ 0\\ 0\end{pmatrix}$ (27b)
@@
$\mathcal{H}_{\text{Ising}}=\begin{pmatrix}0&0&I\end{pmatrix}\left[\;\prod_{\ell=1}^{L}A^{(\ell)}\;\right]\begin{pmatrix}I\\ 0\\ 0\end{pmatrix}.$ (28)
We immediately notice that since the Pauli operators act on single qubits and have dimension $2\times 2$, each matrix has dimension $6\times 6$ and needs at least 3 qubits to be represented. As a consequence, all the matrices and vectors in Eq. (28) gain a new row and/or column in the following way:
@@
$A^{(\ell)}$ $\longrightarrow\begin{pmatrix}I&0&0&0\\ \sqrt{|J|}Z_{\ell}&0&0&0\\ gX_{\ell}&\text{sign}(J)\sqrt{|J|}Z_{\ell}&I&0\\ 0&0&0&I\end{pmatrix}.$ (29c)
@@
Figure 6: Normalization factor $N$ (see Eq. (4)) depending on $g$ and $J$.
Note that to construct a block encoding of $A^{(\ell)}$, we have to normalize it by a factor $N\geq\norm{\mathcal{M}\left[A^{(\ell)}\right]}$, which can be chosen constant for all $\ell$ due to translation invariance. This has the same effect as dividing the Hamiltonian parameters $J$ and $g$ by $N^{L}$:
@@
In Fig. 6, we plot the normalization factor $N$ for different values of $g$ and $J$. We notice that in the limit of vanishing $g$ and $J$, $A^{(\ell)}$ becomes a diagonal matrix with eigenvalues $1$ and 00, so that we get $\left\|\mathcal{M}\left[A^{(\ell)}\right]\right\|\rightarrow 1$.
@@
$A^{(\ell)}$ $\longrightarrow\begin{pmatrix}I&0&0&0\\ \sqrt{|J|}Z_{\ell}&0&0&0\\ gX_{\ell}&\text{sign}(J)\sqrt{|J|}Z_{\ell}&I&0\\ 0&0&0&\zeta^{1/L}I\end{pmatrix}$ (31c)
@@
$J$ $\longrightarrow\frac{J}{2N^{L}},$ (33a)
$g$ $\longrightarrow\frac{g}{2N^{L}}.$ (33b)
@@ VI.2 Heisenberg model
Next, we discuss the Heisenberg model governed by the Hamiltonian:
$\mathcal{H}_{\text{Heis}}$ $=\sum_{\ell=1}^{L-1}\left(J_{X}X_{\ell}X_{\ell+1}+J_{Y}Y_{\ell}Y_{\ell+1}+J_{Z}Z_{\ell}Z_{\ell+1}\right)$
$+\sum_{\ell=1}^{L}\left(g_{X}X_{\ell}+g_{Y}Y_{\ell}+g_{Z}Z_{\ell}\right).$ (34)
@@
$A^{(\ell)}=\begin{pmatrix}I&0&0&0&0\\ \sqrt{|J_{X}|}X_{\ell}&0&0&0&0\\ \sqrt{|J_{Y}|}Y_{\ell}&0&0&0&0\\ \sqrt{|J_{Z}|}Z_{\ell}&0&0&0&0\\ g_{X}X_{\ell}+g_{Y}Y_{\ell}+g_{Z}Z_{\ell}&\text{sign}(J_{X})\sqrt{|J_{X}|}X_{\ell}&\text{sign}(J_{Y})\sqrt{|J_{Y}|}Y_{\ell}&\text{sign}(J_{Z})\sqrt{|J_{Z}|}Z_{\ell}&I\\ \end{pmatrix}.$ (35)
The matrix $A^{({\ell})}$ in Eq. (35) has dimension $10\times 10$ and therefore needs to be expanded into a $16\times 16$ matrix (analogous to the Ising MPO model) to be represented by $4$ qubits. Moreover, another dilation ancillary is added to build $U_{A^{(\ell)}}$ through a QR decomposition.
@@ VI.3 Exponentially decaying XY model
$\mathcal{H}_{\text{XY}_{\gamma}}=\sum_{\ell_{1}=1}^{L}\sum_{\ell_{2}=\ell_{1}+1}^{L}e^{-\gamma|\ell_{1}-\ell_{2}|}\left(J_{X}X_{\ell_{1}}X_{\ell_{2}}+J_{Y}Y_{\ell_{1}}Y_{\ell_{2}}\right)$ (36)
with $\gamma>0$.
If we wanted to represent such a Hamiltonian with a linear combination of unitaries, we would need to consider $M=\mathcal{O}(L^{2})$ terms, which results in a gate decomposition complexity of $\mathcal{O}(L^{2}\cdot\log^{\beta}(L))$ and $\mathcal{O}(\log(L))$ ancillaries, as discussed in section V.
@@
$A^{(\ell)}=\begin{pmatrix}I&0&0&0\\ e^{-\gamma}\sqrt{|J_{X}|}X_{\ell}&e^{-\gamma}I&0&0\\ e^{-\gamma}\sqrt{|J_{Y}|}Y_{\ell}&0&e^{-\gamma}I&0\\ 0&\text{sign}(J_{Y})\sqrt{|J_{X}|}X_{\ell}&\text{sign}(J_{Y})\sqrt{|J_{Y}|}Y_{\ell}&I\\ \end{pmatrix}.$ (37)
Our MPO block encoding technique would then be able to represent this Hamiltonian with a number of one- and two-qubit scaling only linearly with the system size $L$. The number of ancillaries would also scale as $\mathcal{O}(L)$ as well as the cost of the signal processing, giving a total cost for a full QET circuit of only $\mathcal{O}(d\cdot L)$, where $d$ is the degree of the polynomial.
It’s also worth noticing that with LCU we would need to handle exponentially small coefficients, which might cause practical challenges and require a truncation for very small terms. With the MPO block encoding, on the other hand, the factor $e^{-\gamma}$ is the same in every matrix $A^{(\ell)}$, and we can get an exact representation without any truncation.
VI.4 Spinless Fermi-Hubbard
We briefly report also the one-dimensional spinless Fermi-Hubbard as another example of Hamiltonian presenting a constant $\chi$. Unlike the other examples shown so far, this is a fermionic Hamiltonian and needs to be embedded in the qubit system by a fermions-to-qubits mapping, e.g., the Jordan-Wigner mapping [22]. The embedded Fermi-Hubbard Hamiltonian looks then like:
$\mathcal{H}_{\text{FH}}$ $=J\sum_{\ell=1}^{L-1}\left(S^{+}_{\ell}S^{-}_{\ell+1}+S^{-}_{\ell}S^{+}_{\ell+1}\right)$
$+u\sum_{\ell=1}^{L-1}n_{\ell}n_{\ell+1}$ (38)
where the rising and lowering operators are defined as:
$S^{\pm}_{\ell}=\frac{X_{\ell}\mp iY_{\ell}}{2}$ (39)
and the number operator is:
$n_{\ell}=S^{+}_{\ell}S^{-}_{\ell}.$ (40)
The MPO of such a Hamiltonian will then look like:
$A^{(\ell)}=\begin{pmatrix}I&0&0&0&0\\ \sqrt{|J|}S^{+}_{\ell}&0&0&0&0\\ \sqrt{|J|}S^{-}_{\ell}&0&0&0&0\\ \sqrt{|u|}n_{\ell}&0&0&0&0\\ 0&\text{sign}(J)\sqrt{|J|}S^{-}_{\ell}&\text{sign}(J)\sqrt{|J|}S^{+}_{\ell}&\text{sign}(u)\sqrt{|u|}n_{\ell}&I\\ \end{pmatrix}.$ (41)
This model can be further generalized for the spinful case. Still, this simple system suffices to show one possible example of a fermionic Hamiltonian that can be efficiently encoded with our algorithm.
VI.5 Tensor product of sum of Paulis
@@
$\mathcal{H}_{\text{TP}}=\bigotimes_{\ell=1}^{L}\left(\alpha_{\ell}I+\beta_{\ell}X_{\ell}+\gamma_{\ell}Y_{\ell}+\delta_{\ell}Z_{\ell}\right).$ (42)
@@
$A^{(\ell)}=\left(\alpha_{\ell}I+\beta_{\ell}X_{\ell}+\gamma_{\ell}Y_{\ell}+\delta_{\ell}Z_{\ell}\right)$ (43)
@@
$A^{(\ell)}\longrightarrow\begin{pmatrix}\alpha_{\ell}I+\beta_{\ell}X_{\ell}+\gamma_{\ell}Y_{\ell}+\delta_{\ell}Z_{\ell}&0\\ 0&\zeta^{1/L}I\end{pmatrix}$ (44)
@@
$\bigotimes_{\ell=1}^{L}\left(\alpha_{\ell}I+\beta_{\ell}X_{\ell}+\gamma_{\ell}Y_{\ell}+\delta_{\ell}Z_{\ell}\right)+\zeta I=\\ \begin{pmatrix}I&I\end{pmatrix}\left[\prod_{\ell=1}^{L}\begin{pmatrix}\alpha_{\ell}I+\beta_{\ell}X_{\ell}\\ +\gamma_{\ell}Y_{\ell}+\delta_{\ell}Z_{\ell}&0\\ 0&\zeta^{1/L}I\end{pmatrix}\right]\begin{pmatrix}I\\ I\end{pmatrix},$ (45)
@@
VI.6 Eigenstate filtering
Figure 8: 8a Original eigenvalues for tensor product of Pauli sums with $L=3$, $\alpha=\{0.7,1.2,-0.3\}$, $\beta=\{-1,0.4,0.5\}$, $\gamma=\{0,0.3,0.5\}$, $\delta=\{0.1,0,1.2\}$, $\zeta=1.7$ and $N=1.72$. 8b Filtering function (in pink) plus QET-transformed eigenvalues. Phases were calculated via the pyqsp package with degree $d=30$ and threshold $\Delta_{\lambda}=0.1$.
Finally, we test our block encoding algorithm, as well as our signal processing technique, by applying an eigenstate filtering function to the Hamiltonian defined in (42). In particular, we want to perform the following transformation over $\mathcal{H}$ [23]:
$f_{d}(x,\Delta_{\lambda})=\frac{T_{d}\left(-1+2\frac{x^{2}-\Delta_{\lambda}^{2}}{1-\Delta_{\lambda}^{2}}\right)}{T_{d}\left(-1+2\frac{-\Delta_{\lambda}^{2}}{1-\Delta_{\lambda}^{2}}\right)}$ (46)
@@ VI.6 Eigenstate filtering
$\text{poly}(\mathcal{H})=\frac{1}{2}\left(f_{\text{QET}}(\mathcal{H})+f_{\text{QET}}(\mathcal{H})^{\dagger}\right).$ (47)
@@ VII Conclusion
Our block encoding method is generally applicable since typical Hamiltonians of interest can be represented as matrix product operators. Moreover, our block encoding technique’s gate decomposition cost grows only linearly with the system size and quadratically with the virtual bond dimension, offering a potential advantage compared to LCU. In section VI we have shown one specific example of a circuit size advantage in the case of an exponentially decaying XY model.
The signal processing operator can have different implementations, as we have shown in section IV, and it incurs a cost growing linearly or quadratically with the number of ancillaries, depending on the chosen implementation and the presence of one extra auxiliary qubit [19, 14, 15].
As previously mentioned, each matrix encoded into a larger unitary must have a bounded norm. Encoding directly $\mathcal{H}$ in a larger unitary requires $\norm{\mathcal{H}}\leq 1$. In our case, we need to normalize each matrix $\mathcal{M}[A^{(\ell)}]$ individually, causing an equivalent rescaling of the Hamiltonian by the factor $N_{\text{MPO}}$ defined by Eq. (11). The required renormalization is a general issue of block encoding and is not specific to our MPO approach. We remark that the spectral norms of the individual matrices $\mathcal{M}[A^{(\ell)}]$ are typically easier to calculate than $\norm{\mathcal{H}}$. However, $N_{\text{MPO}}$ grows exponentially with $L$, which causes an exponentially small success probability for our method in the asymptotic limit. Nevertheless, in appendix B we have shown how rescaling $\mathcal{H}$ could partially solve this issue when considering finite systems.
Instead, the final success probability for a QET circuit is not directly impacted by the normalization factor, which appears in the polynomial transformation argument, see Eq. (IV). Still, an exponential suppression of the eigenvalues might bring some practical disadvantages. This is currently the biggest limitation of our method, but we have been wondering if a different MPO implementation could help to reduce the norm $\|\mathcal{M}[A^{(\ell)}]\|$, for example considering gauge symmetries of the MPO. We have decided to leave this question for future work.
@@
Moreover, the use of adaptive circuits might further reduce the size and depth of the MPO block encoding, as inspired by [25, 26, 27].